# Self-supervised Learning for Acoustic Few-Shot Classification

### v1

@@ Abstract
Labelled data are limited and self-supervised learning is one of the most important approaches for reducing labelling requirements. While it has been extensively explored in the image domain, it has so far not received the same amount of attention in the acoustic domain. Yet, reducing labelling is a key requirement for many acoustic applications. Specifically in bioacoustic, there are rarely sufficient labels for fully supervised learning available. This has led to the widespread use of acoustic recognisers that have been pre-trained on unrelated data for bioacoustic tasks. We posit that training on the actual task data and combining self-supervised pre-training with few-shot classification is a superior approach that has the ability to deliver high accuracy even when only a few labels are available. To this end, we introduce and evaluate a new architecture that combines CNN-based preprocessing with feature extraction based on state space models (SSMs). This combination is motivated by the fact that CNN-based networks alone struggle to capture temporal information effectively, which is crucial for classifying acoustic signals. SSMs, specifically S4 and Mamba, on the other hand, have been shown to have an excellent ability to capture long-range dependencies in sequence data. We pre-train this architecture using contrastive learning on the actual task data and subsequent fine-tuning with an extremely small amount of labelled data. We evaluate the performance of this proposed architecture for ($n$-shot, $n$-class) classification on standard benchmarks as well as real-world data. Our evaluation shows that it outperforms state-of-the-art architectures on the few-shot classification problem.
@@ I Introduction
One of the most promising approaches for reducing labelling requirements is the use of self-supervised learning[2, 3]. Self-supervision can be used in an initial pre-training phase and the networks obtained can subsequently be fine-tuned with a very small amount of labelled data for specific downstream tasks. We are specifically concerned with classification, where the above approach is the basis of $n$-class, $n$-shot classification in which just $n$ labelled samples per class are used to train a classifier for $n$ classes.
Such $n$-class, $n$-shot classification is specifically relevant in bioacoustic. Here, a typical task is the recognition of threatened species[4]. Since sufficient problem-specific labelled train data is rarely available, particularly for cryptic species, the use of recognisers that have been pre-trained on unrelated data is general practice. For example, mammals are commonly classified with recognisers that have been pre-trained on bird data or even generalised audio data, such as AudioSet[5].
@@
Our architecture is based on a combination of convolutional neural network (CNN) blocks for feature pre-processing with a Structured State Space Sequence model (S4). The use of S4 is motivated by the fact that S4 architectures, including Mamba, are specifically designed to model long sequences and have been proven to have great potential for modelling long-range dependencies in sequence data. SSMs reach equivalent performance to other sequence models, specifically Transformers, while significantly reducing the computational effort[7, 8]. CNNs, on the other hand, are the most commonly used approach for supervised learning of audio data[9] and are effective at feature processing. We thus use a CNN structure based on ResNet[10] for initial feature processing, leaving the time-dimension untouched, and subsequently process the time series of preprocessed features with an S4 architecture. We specifically use S4D[11, 12], an improved version of S4 that reduces the computational requirements.
We evaluate the use of our architecture for five-class five-shot classification on standard benchmark data (ESC 50) as well as on real-world data recorded for 10 different frog species recorded in the field in Queensland, Australia[13]. We pre-train the embedder on the actual problem data and subsequently fine-tune the downstream classifier with labelled samples from the same data set. This truthfully reflects the real application scenario, where unlabelled data from the specific application scenario is generally available in copious amounts while labelling is very expensive so that only a few labelled samples are available. It is important to note that this is subtly different from other versions of the $n$-class, $n$-sample few-shot classification, where data used in the pre-training are different from those used for the fine-tuning [14]. The latter tests the generalisation ability of the network to new unseen classes, whereas we are interested in achieving increased performance by specialising an embedder for a specific application scenario using only a very limited labelling budget.
@@ II-B Contrastive learning method
Motivated by SimCLR[15] and COLA[16], we learn a robust representation of unlabelled audio signals by training our proposed AcousticSSM with a contrastive loss function. We investigated two different contrastive learning methods in our experiments. The first method (AcousticSSM1) assigns high similarity to audio segments extracted from the same audio recording and low similarity to audio segments extracted from different audio recordings. The loss function maximises the agreement between an anchor segment and a related positive segment from the same audio clip while minimising the agreement between this anchor segment and negative segments from unrelated clips. Instead of keeping a memory bank of negatives after picking one anchor segment, we use a bullet strategy: positive segments are defined as negative segments for all other anchors in one batch.
@@ III Experimental Evaluation
We evaluate our method on the well-known Environmental Sound Classification benchmark ESC50[18] as well as on a real-world bioacoustic dataset. To the best of our knowledge, no existing benchmark for the ESC50[18] five-way five-shot classification tasks is available. ESC50 consists of 2,000 5-second samples of environmental recordings equally distributed across 50 classes (40 clips per class). Our bioacoustic dataset contains calls of 5 frog species recorded at 4-Mile-Creek, Townsville, Queensland, Australia in 2020 [13] (L rubella, L rothii, L pornatum, C novae, L caerulea). [^1] Thanks to Lin Schwartzkopf and Slade Allen-Ankins (James Cook University, Townsville) for making this dataset available. We use 50 unlabelled samples of each species and 5 labelled samples for fine-tuning. To fit our 5-class 5-shot problem, we group ESC 50 into 10 groups $G_{1},\ldots,G_{10}$ with each group containing 5 classes and group divisions according to coarse semantic categories (animal vocalisations, human vocalisations, natural environment sounds, interior built environment, urban exterior sounds, …).
@@
In each experiment, one group $G_{i}$ is chosen. For each of the 5 classes in $G_{i}$ 5 labelled samples are reserved for fine-tuning and all other samples (but without labels) are used for pre-training. We use ADAM for pre-training and SGD for fine-tuning. For pre-training, we apply a learning rate of 0.0001 for 500 epochs. For fine-tuning, we apply a learning rate of 0.006 for 50 epochs, where the best parameter was optimised using a parameter sweep.  [^2] The learning rate for fine-tuning was optimised using a parameter sweep between 0.005 and 0.01.

### v2

@@ Abstract
Labelled data are limited and self-supervised learning is one of the most important approaches for reducing labelling requirements. While it has been extensively explored in the image domain, it has so far not received the same amount of attention in the acoustic domain. Yet, reducing labelling is a key requirement for many acoustic applications. Specifically in bioacoustic, there are rarely sufficient labels for fully supervised learning available. This has led to the widespread use of acoustic recognisers that have been pre-trained on unrelated data for bioacoustic tasks. We posit that training on the actual task data and combining self-supervised pre-training with few-shot classification is a superior approach that has the ability to deliver high accuracy even when only a few labels are available. To this end, we introduce and evaluate a new architecture that combines CNN-based preprocessing with feature extraction based on state space models (SSMs). This combination is motivated by the fact that CNN-based networks alone struggle to capture temporal information effectively, which is crucial for classifying acoustic signals. SSMs, specifically S4 and Mamba, on the other hand, have been shown to have an excellent ability to capture long-range dependencies in sequence data. We pre-train this architecture using contrastive learning on the actual task data and subsequent fine-tuning with an extremely small amount of labelled data. We evaluate the performance of this proposed architecture for ($n$-shot, $n$-way) classification on standard benchmarks as well as real-world data. Our evaluation shows that it outperforms state-of-the-art architectures on the few-shot classification problem.
@@ I Introduction
One of the most promising approaches for reducing labelling requirements is the use of self-supervised learning[2, 3]. Self-supervision can be used in an initial pre-training phase and the networks obtained can subsequently be fine-tuned with a very small amount of labelled data for specific downstream tasks. We are specifically concerned with classification, where the above approach is the basis of $n$-way, $n$-shot classification in which just $n$ labelled samples per class are used to train a classifier for $n$ way.
Such $n$-way, $n$-shot classification is specifically relevant in bioacoustic. Here, a typical task is the recognition of threatened species[4]. Since sufficient problem-specific labelled train data is rarely available, particularly for cryptic species, the use of recognisers that have been pre-trained on unrelated data is general practice. For example, mammals are commonly classified with recognisers that have been pre-trained on bird data or even generalised audio data, such as AudioSet[5].
@@
Our architecture is based on a combination of convolutional neural network (CNN) blocks for feature pre-processing with a Structured State Space Sequence model (S4). The use of S4 is motivated by the fact that S4 architectures, including Mamba, are specifically designed to model long sequences and have been proven to have great potential for modelling long-range dependencies in sequence data. SSMs reach equivalent performance to other sequence models, specifically Transformers, while significantly reducing the computational effort[7, 8, 9, 10]. CNNs, on the other hand, are the most commonly used approach for supervised learning of audio data[11] and are effective at feature processing. We thus use a CNN structure based on ResNet[12] for initial feature processing, leaving the time-dimension untouched, and subsequently process the time series of preprocessed features with an S4 architecture. We specifically use S4D[13, 14], an improved version of S4 that reduces the computational requirements.
We evaluate the use of our architecture for five-way five-shot classification on standard benchmark data (ESC 50) as well as on real-world data recorded for 10 different frog species recorded in the field in Queensland, Australia[15]. We pre-train the embedder on the actual problem data and subsequently fine-tune the downstream classifier with labelled samples from the same data set. This truthfully reflects the real application scenario, where unlabelled data from the specific application scenario is generally available in copious amounts while labelling is very expensive so that only a few labelled samples are available. It is important to note that this is subtly different from other versions of the $n$-way, $n$-shot few-shot classification, where data used in the pre-training are different from those used for the fine-tuning [16]. The latter tests the generalisation ability of the network to new unseen classes, whereas we are interested in achieving increased performance by specialising an embedder for a specific application scenario using only a very limited labelling budget.
@@ II-B Contrastive learning method
Motivated by SimCLR[17] and COLA[18], we learn a robust representation of unlabelled audio signals by training our proposed AcousticSSM with a contrastive loss function. We investigated two different contrastive learning methods in our experiments. The first method (AcousticSSM1) assigns high similarity to audio segments extracted from the same audio recording and low similarity to audio segments extracted from different audio clips. The loss function maximises the agreement between an anchor segment and a related positive segment from the same audio clip while minimising the agreement between this anchor segment and negative segments from unrelated clips. Instead of keeping a memory bank of negatives after picking one anchor segment, we use a bullet strategy: positive segments are defined as negative segments for all other anchors in one batch.
@@ III Experimental Evaluation
We evaluate our method on the well-known Environmental Sound Classification benchmark ESC50[20] as well as on a real-world bioacoustic dataset. To the best of our knowledge, no existing benchmark for the ESC50[20] five-way five-shot classification tasks is available. ESC50 consists of 2,000 5-second samples of environmental recordings equally distributed across 50 classes (40 clips per class). Our bioacoustic dataset contains calls of 5 frog species recorded at 4-Mile-Creek, Townsville, Queensland, Australia in 2020 [15] (L rubella, L rothii, L pornatum, C novae, L caerulea). [^1] Thanks to Lin Schwartzkopf and Slade Allen-Ankins (James Cook University, Townsville) for making this dataset available. We use 50 unlabelled samples of each species and 5 labelled samples for fine-tuning. To fit our 5-way 5-shot problem, we group ESC 50 into 10 groups $G_{1},\ldots,G_{10}$ with each group containing 5 classes and group divisions according to coarse semantic categories (animal vocalisations, human vocalisations, natural environment sounds, interior built environment, urban exterior sounds, …).
@@
In each experiment, one group $G_{i}$ is chosen. For each of the 5 ways in $G_{i}$ 5 labelled samples are reserved for fine-tuning and all other samples (but without labels) are used for pre-training. We use ADAM for pre-training and SGD for fine-tuning. For pre-training, we apply a learning rate of 0.0001 for 500 epochs. For fine-tuning, we apply a learning rate of 0.006 for 50 epochs, where the best parameter was optimised using a parameter sweep.  [^2] The learning rate for fine-tuning was optimised using a parameter sweep between 0.005 and 0.01.